### v2

@@ Abstract
We collect and study 15,097 blocks proposed for inclusion in the Ethereum blockchain during 8 minutes on December 3rd, 2024, corresponding to 39 added blocks. These proposed blocks contain 10,793 unique transactions, and 2,380,014 transaction-block pairings, our primary unit of analysis. We find that 20% of user transactions are delayed: although proposed during a bidding cycle, they are not included in the corresponding winning block. Approximately 30% of such delayed transactions are exclusive to a losing builder. We also identify two arbitrage bots trading between decentralized (DEX) and centralized exchanges (CEX), which are responsible for a significant fraction of the value of proposed blocks. By examining their bidding dynamics, we estimate that the implied price at which these bots trade USDC/WETH and USDT/WETH on CEXes is approximately 2.8 basis points better than the contemporaneous Binance price.
@@ 1 Introduction
inlineinlinetodo: inlinebefore submitting, we should make sure that we are consistent with the naming of builders and relaysinlineinlinetodo: inlinechange ”block value” to ”block’s potential revenues” (or somethign else)inlineinlinetodo: inlineMissing: more info on transaction delay. A plot, and also what these delays constitute in terms of feeinlineinlinetodo: inlineMissing: something that gives a general sense of how much variation there is in terms of prob of success and price (then we can jump to explaining this variation)
Blockchains are often described as immutable ledgers because once a transaction is included in the canonical chain it cannot be altered. This static view, however, obscures the probabilistic and competitive process that determines which transactions become immutable. In Ethereum, this process is governed by the interaction of proposers, builders, and relays. The block proposer has the right to add the next block to the chain and is randomly selected among those who have staked at least 32 ETH. Builders are specialized entities that gather transactions and use algorithms to decide which to include and in what order. Relays act as auctioneers: they collect blocks and payments offered to the proposer (bids) from builders, continuously broadcast the current highest bid, and forward the corresponding block to the proposer when requested. [^1] Relays are “trusted” because they must also prevent proposers from tampering with submitted blocks. For more details, see https://docs.flashbots.net/flashbots-mev-boost/relay. Each auction cycle lasts about 12 seconds, after which the winning block is added to the chain and the process begins again.
Within a cycle, builders continually collect new transactions from users and from searchers, automated bots that generate transactions in response to market conditions (such as price movements) or to other pending transactions. They use these new transactions to update and resubmit blocks. Latency is crucial, as proposers are globally distributed and arbitrage opportunities often depend on rapid movements in off-chain financial markets. To mitigate these frictions, multiple relays operate worldwide, and builders typically submit to several at once. As a result, for each block added to the canonical chain, up to 5,000 candidate blocks are constructed, propagated, and ultimately discarded along with their transactions. Each of these discarded blocks represents an alternative, unrealized history. Measuring how much these unrealized histories differ from the one that ultimately appears on-chain is a key research question for a system that supports about USD 5 billion in daily financial exchanges. Yet, until now, data limitations have prevented researchers from addressing it.
In this paper, we overcome this limitation by assembling a novel dataset of proposed blocks submitted between Ethereum blocks 21,322,622 and 21,322,660 (39 winning blocks), roughly from 2:37:35 PM to 2:45:25 PM UTC on December 3, 2024. This dataset contains 15,097 blocks, 10,793 unique transactions (identified by their hash) and 2,380,014 transaction–block pairs, which constitute our main unit of analysis. Crucially, the blocks in our dataset are not a random sample: they are all blocks submitted to a relay called Agnostic. [^2] Founded in 2022, Agnostic relay held about 20% of the relay market in 2023, declining to around 5% during our study period. This introduces two limitations. First, we observe no submitted blocks for 11 bidding cycles because the proposers assigned to those slots were not connected to Agnostic, which therefore did not collect block submissions. Second, it contains proportionally few blocks from Beaverbuild, a major builder during the study period: in bidding cycles for which we have data, Beaverbuild accounted for 40.2% of all blocks submitted to major relays but it contributed only 9.4% of the blocks submitted to Agnostic (and thus included in our dataset). We later discuss how latency may explain Beaverbuild’s limited presence on Agnostic. With these limitations in mind, we can now discuss our main results.
Our main research question is whether the process of block selection matters for users. We find that it does. Approximately 20% of users’ transactions are delayed: they first appear during an auction cycle in some proposed blocks but not in the winning one, delaying their inclusion on-chain by one or more bidding cycles. Strikingly, about 30% of these delayed transactions initially appear as exclusive, i.e., they are present only in blocks submitted by a single builder. We also examine what determines whether a user’s swap executes in some blocks but not others, and at what price it executes. We document substantial variance in both execution success and prices across blocks containing the same transaction.11todo: 1add values Regression results show that part of this variance can be explained by what other transaction is included in a block: swaps in the same direction as the two main arbitrage bots in our data (see the next paragraph) are about 18% less likely to execute when included in a block that also contains transactions from those bots. By contrast, swaps in the opposite direction are about 1% more likely to execute under the same conditions. A similar, though smaller, pattern holds for prices: swaps in the opposite direction of the bots trade at better prices when included in blocks that also contain their transactions, while swaps in the same direction trade at worse prices. Hence, the process of block selection materially affects users, particularly through the timing of transaction execution.
We then turn to the determinants of which block wins. To do so, we exploit a notable feature of our dataset: 2,981 transactions—35% of all unique transactions—originate from just two addresses that almost exclusively perform swaps and together account for roughly 60% of all transaction fees in the blocks where they appear. The first address, which we label “Titan-bot,” submits transactions only to the builder Titan. The second, “Rsync-bot,” submits transactions almost exclusively to a builder called Rsync, although a few of its transactions occasionally appear in blocks built by both Rsync and Titan. In those cases, however, the fee offered to Titan is always lower than the fee offered to Rsync, with an average difference of 18%. [^3] It is common knowledge that Rsync owns and operates several trading bots. Titan, by contrast, claims not to own or operate any bot. Wu et al. (2025) also identify Titan-bot (“Kyle” in their labels) as exclusive to Titan.
Since these two bots generate most of the fees in each block, their competition sheds light on which builder wins the auction. We assume that each bot’s payment to its builder for block inclusion equals its risk-adjusted expected profit. [^4] This assumption is justified if the bot is owned by the builder itself (Pai and Resnick, 2024), or if the builder rebates to the bot an appropriate share of the fee. For example, a rebate may be calculated in a second-price auction style by comparing the block’s value with and without the bot’s transaction. This is the approach taken by Buildernet, a recent open-source builder operating inside a Trusted Execution Environment (see https://buildernet.org/docs/refunds). By observing each bot’s swap volume, on-chain execution price, and payment for inclusion, we infer an implied centralized exchange (CEX) price—that is, the price of the off-chain leg of the arbitrage trade. We then compare this implied price to the contemporaneous Binance price during the bidding cycle. We find that when both bots compete to rebalance either a USDT/WETH or a USDC/WETH pool, the implied CEX price is roughly 2.8 basis points more favorable than Binance’s contemporaneous price for both bots.
@@ 1.1 Relevant literature
Ethereum’s process of block creation, submission, and selection— known as Proposer-Builder Separation (PBS) — has been the subject of considerable debate (for a summary of this debate, with a particular emphasis on the economic aspects, see John et al. (2025)). Our results contribute directly to these discussions.
@@
One of the main criticism of PBS is that it enables searchers to efficiently extract value from users. Quantitatively, the main concern is arbitrageurs profiting at the expense of on-chain liquidity providers by exploiting price discrepancies between decentralized exchanges (DEXes) and centralized exchanges (CEXes). The resulting loss to liquidity providers is referred to as Loss-vs-Rebalancing (LVR), a concept introduced by Milionis et al. (2022). Subsequent studies have attempted to identify arbitrageurs (Heimbach et al., 2024, Wu et al. (2025)) and quantify LVR (Canidio and Fritsch, 2023; Fritsch and Canidio, 2024). These analyses, however, rely on strong assumptions. This is because on-chain data alone reveals only that a swap occurred, not whether it was part of an arbitrage strategy. Furthermore, the prices at which arbitrageurs trade on centralized venues are usually unobservable. Our dataset overcomes these limitations in two ways. First, we can identify arbitrage bots by observing competition between them for the same arbitrage opportunities. In this respect, our approach resembles that of Aquilina et al. (2022), who identify high-frequency trading races in traditional financial markets by analyzing failed transactions. Their rationale is that when multiple arbitrageurs try to exploit the same arbitrage opportunity, only one transaction will succeed, and all others will fail. Similarly, we observe multiple arbitrage bots trying to exploit the same opportunity by submitting transactions, with only one of them eventually successfully included on-chain. Second, by focusing on builder-integrated searchers engaging in CEX-DEX arbitrage, we can infer their expected profits and thus the price at which they trade on the centralized leg.
The dynamics of builder competition and block inclusion on the Ethereum blockchain have been studied extensively. However, to the best of our knowledge, we are the first to systematically collect data on transactions contained in non-winning blocks. Yang et al. (2024) is the only other paper that analyzes the content of non-winning blocks. They examine competition among builders by comparing builders’ bids to the blocks’ revenues. Several other papers explore different aspects of builder competition. For example, Wu et al. (2024a) and Wu et al. (2024b) study builders’ bidding strategies, both theoretically and through simulation. Bahrani et al. (2024) analyze builder competition theoretically and derive conditions under which the builder and proposer markets become concentrated in equilibrium.
@@
Similarly to our analysis of the competition between Rsync-bot and Titan-bot, Capponi et al. (2024) also studies searchers competing to exploit CEX-DEX arbitrage opportunities. However, they consider an earlier period in which searchers competed by submitting transactions to the public mempool and engaging in priority fee auctions — i.e., because all transactions are included in the block and ordered by priority fee, the winning searcher is the one paying the highest fee. The current market structure and nature of the competition between searchers are very different, with searchers submitting directly to builders, builders selecting which searchers’ transactions to include, and frequent integration between searchers and builders. [^5] See also Capponi et al. (2023), who develop a theory of searchers’ competition when transactions are sent privately to builders.
Also related are Öz et al. (2023) and Schwarz-Schilling et al. (2023), who study how proposers decide when to request the winning block. They argue that this choice is partially strategic, as builders’ bids tend to increase during the auction cycle. Delaying the selection of the winning block can, therefore, lead to higher payments to the proposer. Wahrstätter et al. (2023) and Heimbach et al. (2023) provide an empirical analysis of Ethereum block production from September 2022 to May 2023, a period during which proposer-builder separation (PBS) became the dominant mechanism for block construction on Ethereum. Finally, among the studies dealing with the relationship between searchers and builders, the most relevant ones are Pai and Resnick (2024) and Wu et al. (2025). Pai and Resnick (2024) develops a theory of integration between builders and searchers based on the idea that an integrated searcher’s payment to a builder reflects the full value of including the transaction. Wu et al. (2025) uses on-chain data to identify searchers performing CEX-DEX arbitrage, and study their relationship with the major builders.
@@
Finally, in our companion paper Pahari and Canidio (2025), we use the same data to measure the share of transactions that are exclusive to a builder, meaning they appeared only in blocks proposed by the winning builder during the auction cycle. This question is particularly important given the high concentration in the builders’ market: the top builder (Titan) produces about 50% of all blocks, while the top three account for more than 90%. [^6] For up-to-date statistics, see https://www.relayscan.io/. Buildernet is a joint project by Flashbots (who operated an independent builder until mid-2025) and Beaverbuild (who both participate in Buildernet and run their own builder). Such concentration gives a few entities disproportionate influence over which transactions are included, undermining Ethereum’s goals of decentralization and permissionlessness. We find that between 77% and 84% of the total value of transactions included on-chain is attributable to exclusive transactions, confirming that exclusivity is the main driver of builders’ profits and of PBS auction outcomes.
@@
The remainder of the paper is organized as follows. The next section describes how we constructed our dataset and provides additional summary statistics. The following section provides an in-depth analysis of the auction cycle leading to the addition of Block 21322649 to the Ethereum blockchain. We then extend our analysis to all auction cycles in our dataset. The last section concludes.

### v3

@@ Abstract
Blockchain’s economic value lies in enabling financial and economic transactions that do not require trusted, centralized intermediaries. In practice, however, transactions must pass through several intermediaries before being included on-chain. We study empirically whether this process undermines blockchain’s stated benefits by assembling a novel dataset of 15,097 non-winning Ethereum blocks—blocks proposed by builders but not ultimately selected for inclusion. We show that 21% of user transactions are delayed: although proposed in some candidate blocks, they are not included in the winning block. Approximately 30% of these delayed transactions are exclusive to a single losing builder, indicating that transaction routing materially affect inclusion outcomes. We further document substantial heterogeneity in execution quality: both the probability of successful execution and the execution price of users’ swaps vary across candidate blocks. Finally, we study two arbitrage bots trading between decentralized (DEX) and centralized exchanges (CEX). We document intense competition for the same arbitrage opportunities and estimate that these bots trade USDC/WETH and USDT/WETH on centralized exchanges at prices approximately 2.8 basis points more favorable than contemporaneous Binance prices.
Keywords: Ethereum, Proposer-Builder Separation (PBS), Order Flow.
@@ 1 Introduction
Blockchains are often described as “immutable ledgers” because, once a transaction is included in the canonical chain, it triggers a deterministic sequence of irreversible state changes. This immutability is central to blockchain’s economic value, as it enables financial and economic transactions that do not rely on trusted, centralized intermediaries.
This static view, however, obscures the probabilistic and competitive process that determines which transactions become immutable. In Ethereum, transactions are generated either by users or by searchers—automated bots that create transactions in response to, for example, price movements in financial markets. Users and searchers submit transactions to builders through several channels: the public mempool (a publicly observable repository of pending transactions), direct bilateral communication, or intermediaries known as private mempools (or private RPCs). Builders aggregate and order these transactions into candidate blocks and submit them to relays, each of which forwards a single block to the proposer, who ultimately appends the next block to the chain. [^1] See Section 1.1 for a detailed description of this process, commonly referred to as Proposer–Builder Separation (PBS). The overall process resembles an auction aimed at selecting the block containing the most valuable transactions, as determined by their fees. This auction, however, is highly noisy: the value of including a transaction often depends on rapid price movements in financial markets; builders continuously construct and resubmit blocks; and communication latency plays a central role in determining the final outcome. As a result, for each block ultimately added to the canonical chain, up to 5,000 candidate blocks are constructed, propagated, and then discarded together with their transactions.
An important research question is therefore whether the intermediaries that control transaction inclusion undermine blockchain’s promise of disintermediation. Addressing this question empirically requires comparing discarded Ethereum blocks both with one another and with the block that ultimately appears on-chain. If, for example, an intermediary’s behavior causes a user transaction to appear only in a subset of candidate blocks, access to the blockchain may be delayed or denied because the winning block may omit that transaction. More subtly, transaction outcomes—particularly for swaps—are far less deterministic once we recognize that the same transaction may be included differently across candidate blocks, either by occupying different positions or by being bundled with different other transactions. At the same time, it is also possible that the process of block construction and selection has little impact on ordinary users if all proposed blocks are sufficiently similar from their perspective. Despite the importance of this question, empirical analysis has until recently been infeasible because discarded blocks are not publicly disclosed.
In this paper, we overcome this limitation by assembling novel dataset of non-winning blocks submitted between Ethereum blocks 21,322,622 and 21,322,660 (39 winning blocks), roughly from 2:37:35 PM to 2:45:25 PM UTC on December 3, 2024. The dataset is approximately 20 GB in size and contains 15,097 blocks, 10,793 unique transactions (identified by their hashes), and 2,380,014 transaction–block pairs. [^2] The dataset includes all blocks submitted during the study period through a relay called Agnostic. See Section 2 for a detailed description of the data. We simulate each candidate block to recover transaction-level execution outcomes, and we complement these data with several publicly available sources, including second-by-second price data from Binance.
Our main finding is that the behavior of intermediaries controlling transaction inclusion generates widespread delays. Approximately 21% of users’ transactions appear in some proposed blocks but not in the corresponding winning block, postponing on-chain inclusion by one or more blocks. [^3] We distinguish between users and searchers (see Section 2.2 for a description of the heuristic used to make this distinction). Searchers’ transactions are almost never delayed: they are either included immediately or they are discarded. Strikingly, about 30% of these delayed user transactions appear exclusively in blocks submitted by a single losing builder. This pattern is difficult to rationalize, as users do not benefit from sharing transactions exclusively with a single builder rather than privately with multiple builders. We also identify 12 user transactions that remain exclusive to a losing builder for two or three consecutive auctions before later appearing in the public mempool. These transactions are not only delayed but also subsequently exposed to potential attacks; one is in fact subject to a sandwich attack. In four cases, the same builder that initially had exclusive access ultimately includes the transaction on-chain. We speculate that a private mempool operator initially shares transactions exclusively with a single builder and reverts to the public mempool when that builder fails to win successive auctions.
We also find that the process of block construction and selection introduces substantial randomness into execution outcomes, with execution risk concentrated among a small subset of transactions. To establish this result, we consider each user swap in our data and all blocks in which such swap appears. We simulate those blocks to determine the swap’s execution had that block been included on-chain. We then compute the fraction of blocks in which the swap executes successfully, and average and standard deviation of its execution prices across all blocks in which it appears (conditional on successful execution). We find that the vast majority of users’ swaps (96.7%) execute successfully in all blocks in which they appear. However, a small fraction fail very frequently: 1.7% fail in at least half of the blocks in which they appear, and 1.1% fail in more than 80% of such blocks. Execution prices are similarly skewed. For half of the swaps in our sample, execution prices are quite concentrated: at least two-thirds of realizations deviate by less than 4.3 basis points from the swap’s average execution price across blocks. By contrast, for 7% of swaps execution prices are highly dispersed, with at least one-third of realizations deviating by more than 100 basis points from the swap’s average.
Regression results show that a nontrivial share of the variation in both execution success and execution prices can be explained by the presence of other transactions in the same block. Restricting the sample to swaps that sometimes fail and including transaction fixed effects, we find that swaps trading the same asset pair and in the same direction as the two main arbitrage bots in our data (see the next paragraph) are approximately 18% less likely to execute when included in a block that also contains transactions from those bots. By contrast, swaps trading the same asset pair but in the opposite direction are about 1% more likely to execute under the same conditions. A similar, though quantitatively smaller, pattern emerges for prices: swaps in the opposite direction of the bots trade at more favorable prices when included in blocks that also contain bot transactions, whereas swaps in the same direction trade at worse prices.
We then turn to the determinants of which block wins. To do so, we exploit a salient feature of our dataset: roughly one-third of all transactions originate from just two addresses that almost exclusively perform swaps. We identify these addresses as competing arbitrage bots. They frequently trade the same asset pairs in the same direction; they frequently replace their transactions within an auction cycle; and, when they compete for the same arbitrage opportunity, their swaps are mutually exclusive, in the sense that no proposed block contains transactions from both bots. These bots’ transactions are also highly valuable, accounting on average for about 60% of the total transaction fees in the blocks in which they appear. Their competition therefore plays a central role in determining auction outcomes.
We study this competition by simulating each block that contains a swap transaction from either bot and extracting the associated swap volume, on-chain execution price, and transaction fee. If we further assume that each bot’s fee equals its expected profit from inclusion on-chain, we can infer an implied centralized exchange (CEX) price—that is, the price of the off-chain leg of the arbitrage trade—at different points in time, corresponding to the block construction timestamp. [^4] This assumption is justified by the fact that each bot appears to be integrated with or exclusive to a different builder (see Section 2.2). If the assumption is relaxed, the bot’s fee represents only a lower bound on expected profits, implying that the implied CEX prices we recover may be even more favorable to the bots than our estimates suggest. We find that when the bots compete to rebalance either a USDT/WETH or a USDC/WETH pool, their implied CEX prices are nearly identical, indicating intense competition. Moreover, these implied prices are approximately 2.8 basis points better than contemporaneous Binance prices for both bots.
1.1 Background: the journey of an Ethereum transaction.
In Ethereum, the right to add the next block to the blockchain is assigned to a proposer, randomly selected from validators who have staked at least 32 ETH. In the vast majority of cases (over 95%), the proposer delegates block construction to professional entities known as builders. Builders collect transactions from both the public mempool and private channels and use sophisticated algorithms to determine which transactions to include in a block and in what order. This arrangement is known as Proposer–Builder Separation (PBS).
The relation between builders and proposers is facilitated by relays. Their primary role is to prevent proposers from altering the submitted blocks. [^5] Technically, this is achieved by requiring proposers to sign an empty block, which is then completed by the relay. For more details, see https://docs.flashbots.net/flashbots-mev-boost/relay. Once a builder creates a block, it submits it to a relay along with a proposed payment to the proposer. The relay then forwards the highest-paying block and its corresponding payment to the proposer upon request. This process resembles an ascending price auction because the relay continuously broadcasts the value of the highest-paying bid—though this broadcast is observed with a delay due to network latency. An auction cycle lasts approximately 12 seconds, after which the winning block is selected, and a new auction begins. As already mentioned, latency plays a crucial role, as proposers are geographically dispersed, and arbitrage transactions often depend on rapid price movements in traditional “off-chain” financial markets. To account for geographic dispersion, multiple relays operate worldwide, with builders submitting to several at once. Our dataset, for instance, comes from a relay called Agnostic Relay.
Users can submit transactions through several channels. The default option is to broadcast them to the public mempool, a publicly observable repository of pending transactions maintained collectively by Ethereum nodes. Broadcasting transactions to the public mempool, however, exposes users to attacks. For example, when a user submits a swap on an automated market maker (AMM), a malicious searcher may mount a sandwich attack by front-running the swap with the same trade and subsequently back-running it with the opposite trade, thereby “buying low” and “selling high” at the user’s expense. To mitigate this risk, many users rely on private mempools—services that allow transactions to be sent directly to one or more builders without being broadcast to the public mempool. During the study period, five major private mempool operators were active. Four of them—MEV Blocker, Flashbots Protect, Merkel, and Blink—state in their documentation that they privately share transactions with multiple builders (see janicot2025privatemevprotectionrpcs for a detailed comparison). By contrast, the fifth operator, MetaMask Smart Transactions, does not disclose its transaction-sharing policy.
The choice between submitting a transaction via the public mempool or through one of the private mempools is typically determined by the user’s wallet, a software application that automates transaction creation and broadcasting. Wallets generally provide a default option that users can modify—although this is not always the case, especially for mobile wallets. For example, by default, users of the popular MetaMask wallet submit Ethereum transactions via MetaMask Smart Transactions.
@@ 1.2 Relevant literature
Ethereum’s process of block creation, submission, and selection has been the subject of considerable debate (for a summary of this debate, with a particular emphasis on the economic aspects, see JOHN2025102718). Our results contribute directly to these discussions.
@@
One of the main criticism of PBS is that it enables searchers to efficiently extract value from users. Quantitatively, the main concern is arbitrageurs profiting from price discrepancies between decentralized exchanges (DEXes) and centralized exchanges (CEXes). The resulting loss to on-chain liquidity providers is referred to as Loss-vs-Rebalancing (LVR), a concept introduced by milionis2022automated. Subsequent studies have attempted to identify arbitrageurs (heimbach2024non, wu2025measuring) and quantify LVR (canidio2023arbitrageurs; fritsch2024measuring). These analyses, however, must rely on strong assumptions because on-chain data alone do not reveals whether a given swap was part of an arbitrage strategy. Furthermore, the prices at which arbitrageurs trade on centralized venues are usually unobservable.
Our dataset overcomes these limitations in two ways. First, we can identify arbitrage bots by observing their competition for the same arbitrage opportunities. In this respect, our approach resembles that of aquilina2022quantifying, who identify high-frequency trading races in traditional financial markets by analyzing all messages received by the London Stock Exchange over a 9 week period. Their rationale is that when multiple arbitrageurs try to exploit the same arbitrage opportunity, only one transaction will succeed, and all others will fail. Similarly, we observe multiple arbitrage bots trying to exploit the same opportunity by submitting transactions, with only one of them eventually successfully included on-chain. Second, by focusing on builder-integrated searchers engaging in CEX-DEX arbitrage, we can infer their expected profits and thus the price at which they trade on centralized exchanges.
An important step in our analysis is establishing that our two main arbitrage bots are integrated with, or exclusive to, a different builder—one with a builder called Titan and the other with a builder called Rsync. Our analysis therefore relates to existing studies of the relationship between searchers and builders, in particular pai2024structural and wu2025measuring. pai2024structural develop a theory of vertical integration between builders and searchers, in which an integrated searcher’s payment to the builder reflects the full value of including its transaction. wu2025measuring use on-chain data to identify searchers engaging in CEX–DEX arbitrage and study their interactions with major builders; they also identify the two bots we analyze as integrated with, or exclusive to, Titan and Rsync, respectively.
@@
capponi2024price also studies the competition between arbitrage bots for the same CEX-DEX arbitrage opportunities. However, they consider an earlier period in which searchers competed by submitting transactions to the public mempool and engaging in priority fee auctions — i.e., because all transactions are included in the block and ordered by priority fee, the winning searcher is the one paying the highest fee. The current market structure and nature of the competition between searchers are very different, with searchers submitting directly to builders, builders selecting which searchers’ transactions to include, and frequent integration between searchers and builders. [^6] See also capponi2024price-private, who develop a theory of searchers’ competition when transactions are sent privately to builders.
Existing work studying the strategic behavior of intermediaries responsible for transaction inclusion has largely focused on theoretical analysis. lehar2023battle and CAPPONI2025104132 study the trade-off between submitting transactions via the public mempool and via private channels, and characterize the resulting equilibrium sorting of transactions across submission routes; CAPPONI2025104132 also provides empirical support for its model using data from Flashbots Protect. halaburda2025impossibility develop a theoretical model of transaction inclusion in blockchains that explicitly incorporates competition among block builders and intermediaries responsible for routing users’ transactions. They show that equilibrium outcomes are centralized, with in one or a few entities controlling transactions inclusion. We complement these theoretical arguments by providing empirical evidence on how these intermediaries affect Ethereum users.
@@
In our companion paper pahari2025exclusiveethereumtransactionsevidence, we use the same data to measure the share of transactions that are exclusive to a builder, meaning they appeared only in blocks proposed by the winning builder during the auction cycle. This question is particularly important given the high concentration in the builders’ market: the top builder (Titan) produces about 50% of all blocks, while the top three account for more than 90%. [^7] For up-to-date statistics, see https://www.relayscan.io/. Buildernet is a joint project by Flashbots (who operated an independent builder until mid-2025) and Beaverbuild (who both participate in Buildernet and run their own builder). We find that between 77% and 84% of the total value of transactions included on-chain is attributable to exclusive transactions, confirming that exclusivity is the main driver of builders’ profits and of PBS auction outcomes.
Finally, to the best of our knowledge, yang2025decentralization is the only other paper that examines the content of non-winning blocks. They study builder competition by comparing bids with block revenues (i.e., total fees paid by the transactions included in a proposed block), but do not analyze individual transactions in non-winning blocks.
@@
The remainder of the paper is organized as follows. The next section describes our dataset and provides additional summary statistics. The following section provides an in-depth analysis of the auction cycle leading to the addition of Block 21322649 to the Ethereum blockchain. We then extend our analysis to all auction cycles in our dataset. The last section concludes.